Suppose that the utility function for two commodities is:

U (q1, q2) = q1α q2(1-α)

Let the prices of the two commodities be p1 and p2 and let the consumer’s income be M.

(1) Check the properties of marginal utilities. In particular, check whether it satisfies diminishing marginal utilities.

(2) Assuming all income is spent on these two commodities, derive the demand curves for the two commodities.

(3) What happens if U (q1, q2) = q1α q2β?

1
Expert's answer
2021-08-02T14:57:22-0400

Solution


"U_1= aq_1^{a-1}q_2^{1-a}=a\frac{U(q_1,q_2}{q_2}"


"U_2=(1-a)q_1q_2^{-a}=(1-a)\frac{U(q_1,q_2)}{q_2}"


"MRS=\frac{dq_2}{dq_1}=\frac{U_1}{U_2}=\frac{aq_2}{1-aq_1}"


The MRS is positive hence satisfies diminishing marginal utilities


The demand curve





"aq_1(\frac{I-P_1q_1}{P_2})^{1-a}+(1-a)q_1(\frac{I-P_1q_1}{P_2})^a(\frac{-P_1}{P_2})=0"


"q_1(P_1,P_2,I)=\frac{aI}{P_1}"


"q_2(P_1,P_2,I)=\frac{(1-a)I}{P_2}"


"3.) \frac{∂q}{∂p} + \frac{q∂q}{∂y} = \frac{βq}{p }+ q(\frac{αq}{y}) = (\frac{q}{p})(β +\frac{αpq}{y})"


Shows the substitution effect

when "U (q1, q2) = q_1^\u03b1 q_2^\u03b2"